# Minimum-Action Learning: Energy-Constrained Symbolic Model Selection for Physical Law Identification from Noisy Data

### 16 Mar 2026 v1

@@ Results
Architectural sparsification. The $E_{\mathrm{min}}$-driven soft-to-discrete transition produces near-one-hot gate distributions: gate concentration $C_{\mathrm{gate}}=0.99\pm 0.02$ (Herfindahl–Hirschman Index on effective gate contributions $A_{i}|\theta_{i}|$; $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis) for MAL-trained models ($N=10$ seeds) vs. $C_{\mathrm{gate}}=0.14\pm 0.04$ for teacher-forcing-only baselines (Mann–Whitney $U=100$, $p<10^{-4}$; permutation test $p<0.001$, n=10,000n=10{,}000; SI Appendix, Fig. S6). This architectural sparsification is consistent with Clune et al.’s [8] finding that minimizing wiring costs—a proxy for metabolic expenditure—produces modular structure in evolved networks.
Structural parallels across domains. We note intriguing structural parallels between MAL’s modularity emergence and patterns in biological systems, while emphasizing that these are currently analogies rather than formal equivalences:
(i) Perinatal brain development: Neonatal heart rate and breathing movement coordination during quiet (low-energy) vs. active (high-energy) sleep states exhibit integer-ratio (3:1) phase relationships [21, 33], reflecting dynamic coordination of coupled physiological oscillators under energy constraints. This frequency-locking behavior is a manifestation of the metastable coordination dynamics described by Kelso [22, 38, 24], and parallels the nonlinear frequency-locking of baroreceptor and sympathetic rhythms observed in autonomic neural circuits [16].
(ii) Ecological networks: Microbial gene co-expression networks in the TARA Oceans dataset [36] exhibit modular architectures consistent with wiring-cost minimization [8], though the connection to MAL’s training dynamics remains analogical (SI Appendix, Section S6).
We emphasize that these cross-domain parallels are structural analogies, not established causal connections. The training schedule parameters $(\alpha_{E},\tau)$ are externally imposed hyperparameters, not emergent oscillators; and the biological systems involve genuine coupled-oscillator dynamics. Whether energy-constrained optimization generically produces similar architectural motifs across these domains is a hypothesis warranting formal investigation.
Comparison to Alternative Approaches
@@
We benchmarked MAL against five alternatives on the Kepler-with-noise problem (SI Appendix, Section S4, SI Tables S5–S6):
@@
(1) SINDy variants [3, 20, 13]: Vanilla SINDy with naive differentiation (stride $s=1$) fails completely (noise-dominated). However, when equipped with the same wide-stencil preprocessing ($s=10$) used in MAL, vanilla SINDy identifies $r^{-2}$ in 10/10 seeds, GP-SINDy in 8/10, and ensemble-SINDy in 10/10 (SI Table S5). The wide-stencil technique is the critical enabler, not the learning algorithm. SINDy’s advantage is speed (<1<1 s vs. $\sim$835 s); its limitation is the absence of dynamical validation—SINDy returns sparse coefficients but cannot roll out trajectories or verify energy conservation. MAL’s Noetherian diagnostic provides this additional layer.
@@
(2) HNN [18] and LNN [10]: We trained both on identical data (SI Table S6). HNN (17K parameters, 113 s) achieved excellent energy conservation ($\sigma_{H}=4.1\times 10^{-4}$) but learns a black-box Hamiltonian with no interpretable symbolic form. LNN (17K parameters, 242 s) suffered from Hessian singularities, with validation loss diverging to $10^{6}$ throughout training—a known failure mode for LNNs on noisy data where the mass matrix becomes ill-conditioned. Neither method performs basis selection or yields a symbolic force law.
@@
(3) Mathematical LLM (Qwen2-Math 7B): 0% accuracy on inverse problems, 61% on forward derivation—mathematical manipulation without physical intuition (SI Appendix, Section S1).
@@
(4) Physics-informed NN [32]: 85% accuracy when the inverse-square law is pre-specified, but cannot identify the law de novo.
@@
Summary of niche. MAL occupies a distinct position: it combines interpretable symbolic basis selection (shared with SINDy [3]) with dynamical rollout and Noetherian validation (shared with HNN [18]), while adding explicit energy-efficiency optimization. HNNs/LNNs [18, 10] guarantee conservation but produce black-box energy functions; SINDy [3] yields symbolic expressions but without dynamical consistency checks; MAL provides both within a metabolically constrained framework.
Discussion
Key Findings
Our results establish four key findings:
(1) Metabolic constraints improve physical law identification. Embedding energy minimization ($E_{\min}$) alongside information maximization ($I_{\max}$) in the Triple-Action training objective constrains the model selection problem from an unconstrained search to a tractable optimization. The $E_{\mathrm{min}}$ subsystem drives gate crystallization (basis selection) that does not occur under prediction error alone, while reducing training energy by 40%. This parallels Noble’s biological relativity [28]: information content and metabolic cost jointly determine which model structure the network selects.
@@ Discussion
(3) Integer-ratio coincidences suggest cross-scale connections. We observe that major gate transitions coincide with epochs where the designed schedule ratio $\alpha_{E}/\tau$ passes through integer values. We emphasize that this observation is currently correlational: $\alpha_{E}$ and $\tau$ are externally imposed hyperparameters, not emergent oscillators, and the coincidence may reflect the density of rationals rather than dynamical phase-locking. However, the structural parallel to integer-ratio coordination in physiological oscillators [21, 16] and the broader framework of metastable coordination dynamics [34, 22, 38, 23, 24] is intriguing and motivates formal investigation. Recent work on phase-oscillator networks with synaptic plasticity [6] suggests that frequency-locking phenomena may be generic in systems where coupling weights co-evolve with oscillator dynamics—a structural analogue to MAL’s co-evolution of gate logits and force coefficients. A rigorous test would require showing that the crystallization dynamics are sensitive to the schedule trajectory’s passage through integer ratios—e.g., via circle-map analysis of the coupled $\alpha_{E}$-$\tau$ dynamics—which we leave to future work.
@@
(4) Architectural sparsification as a consequence of energy minimization. Clune et al. [8] demonstrated that networks evolving under pressure to minimize connection costs spontaneously develop modular architectures. Our work shows a related pattern in physics model selection: MAL’s sparse gate selection ($C_{\mathrm{gate}}=0.99$) emerges from $\mathcal{L}_{E_{\min}}$, not from a pre-programmed bias. This is consistent with the hypothesis that metabolic constraints on wiring costs drive architectural sparsification across both biological [4] and artificial systems [8].
Implications for AI and Theoretical Biology
@@
Energy-efficient AI. MAL’s 40% energy reduction over prediction-error-only baselines demonstrates that explicit metabolic constraints ($E_{\min}$ optimization) can improve computational efficiency without sacrificing task performance. The mechanism is architectural: $\mathcal{L}_{E_{\min}}$ drives early gate crystallization, reducing the effective parameter count and shortening convergence time. This principle—that sparsity-inducing regularization simultaneously improves interpretability and efficiency—may generalize beyond the current benchmarks to larger-scale scientific discovery tasks.
@@
Energy-conservation-based model selection. A persistent challenge in AI interpretability is distinguishing genuine causal understanding from correlation [30]. The energy-conservation criterion—preferring models that preserve Hamiltonian dynamics over long-horizon rollouts—provides a physics-grounded approach complementary to recent work on learning conserved quantities [41, 37]: models that violate energy conservation reveal their failure to capture true causal structure, even when matching short-term predictions. This criterion is inspired by Noether’s theorem (time-translation symmetry implies energy conservation), though our implementation is a specific instantiation rather than a general Noether construction.
Developmental origins of health and disease. The structural analogy between MAL’s two-phase optimization and brain development suggests testable hypotheses: Does disruption of metabolic coordination (e.g., maternal inflammation [15]) alter the integer-ratio structure of physiological coupling? Could this predict neurodevelopmental outcomes? While the correspondence is currently analogical, the intrinsic crystallization timescales observed in MAL motivate quantitative investigation of equivalent timescales in biological development.
Limitations and Future Directions
@@
Fixed basis library. MAL requires pre-specifying candidate force laws $\{\phi_{i}\}$, with the correct answer included. This makes MAL a model selection framework, not an open-ended discovery system. Basis library sensitivity experiments (SI Appendix, Section S4F) quantify this limitation: adding near-confounders ($r^{-2.5}$, $r^{-1.5}$) reduces the correct-basis rate from 50% to 20%, while removing the correct basis ($r^{-2}$) causes the system to split between alternatives. Distant additions ($r^{2}$, $r^{-4}$, $\ln r$) have no effect. In all cases, the energy-conservation diagnostic remains informative, and uniformly elevated $\sigma_{H}$ across seeds flags potential library inadequacy. Extending to genuinely autonomous discovery—where the algorithm constructs novel functional forms not in the library—requires integration with open-ended symbolic regression methods [40, 11] or LLM-guided symbolic search, which is an important direction for future work.
@@
Cross-scale analogies. The structural parallels between MAL’s training dynamics, physiological coordination, and ocean genomic modularity are currently analogies, not formal equivalences. The training schedule parameters are externally imposed, not emergent oscillators. Establishing formal connections would require, e.g., circle-map analysis showing sensitivity of crystallization to integer-ratio passages, or demonstrating Farey coordination in systems where the schedule is itself learned rather than designed.
Conclusion
We have demonstrated that minimum-action learning—neural network training guided by a Triple-Action functional combining information maximization, energy minimization, and symmetry enforcement—enables energy-constrained identification of physical force laws from noisy data. Validated on two benchmarks (Kepler gravity and Hooke’s linear restoring force), the key technical contributions are: (1) the wide-stencil noise reduction technique that transforms an intractable problem (SNR $\sim 0.02$) into a learnable one (SNR $\sim 1.6$); (2) the soft-to-discrete gate sharpening mechanism that achieves basis selection through energy-driven architectural crystallization, with a physics-informed initialization achieving 10/10 correct selection; and (3) the Noetherian model selection criterion that discriminates correct from incorrect physics via long-horizon energy conservation. Direct comparison against SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: interpretable, energy-constrained model selection combining symbolic basis identification with dynamical validation.
The modular architectures that emerge from this process parallel patterns observed in biological systems under metabolic constraints—including physiological coordination [21] and ocean genomic modularity [36, 8]—though these parallels are currently structural analogies rather than formal equivalences. Establishing whether a unified mathematical framework underlies self-organization across these domains is an important open question.
Future work will extend MAL to non-central forces, dissipative systems, and real observational data, investigate whether the training schedule’s integer-ratio coincidences reflect genuine dynamical phase-locking, and explore whether explicit metabolic constraints improve scientific discovery in more complex domains. As Macklem emphasized [27], physiology must address the spontaneous emergence of self-organized order; our results suggest that embedding biological energy-minimization principles in artificial systems is a productive path toward this goal.
Methods
Data Generation
@@ Methods
Training Protocol
Two-phase schedule (BGNO):
@@
Code and Data Availability

### 25 Mar 2026 v2

@@ II.5 Energy-conservation-based model selection
Architectural sparsification. The $E_{\mathrm{min}}$-driven soft-to-discrete transition produces near-one-hot gate distributions: gate concentration $C_{\mathrm{gate}}=0.99\pm 0.02$ (Herfindahl–Hirschman Index on effective gate contributions $A_{i}|\theta_{i}|$; $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis) for MAL-trained models ($N=10$ seeds) vs. $C_{\mathrm{gate}}=0.14\pm 0.04$ for teacher-forcing-only baselines (Mann–Whitney $U=100$, $p<10^{-4}$; permutation test $p<0.001$, n=10,000n=10{,}000; Supplemental Material, Fig. S6). This architectural sparsification is consistent with Clune et al.’s [11] finding that minimizing wiring costs—a proxy for metabolic expenditure—produces modular structure in evolved networks.
Structural parallels. The sparse architectures produced by MAL’s energy-constrained optimization parallel modular structures observed in networks evolved under wiring-cost pressure [11, 12]. Whether energy-constrained optimization generically produces similar architectural motifs across biological and artificial systems is a hypothesis warranting formal investigation; we discuss cross-domain parallels further in the Supplemental Material (Sections S5–S6).
II.6 Comparison to alternative approaches
@@
We benchmarked MAL against five alternatives on the Kepler-with-noise problem (Supplemental Material, Section S4, Supplemental Tables S5–S6):
@@
(1) SINDy variants [3, 14, 15]: Vanilla SINDy with naive differentiation (stride $s=1$) fails completely (noise-dominated). However, when equipped with the same wide-stencil preprocessing ($s=10$) used in MAL, vanilla SINDy identifies $r^{-2}$ in 10/10 seeds, GP-SINDy in 8/10, and ensemble-SINDy in 10/10 (Supplemental Table S5). The wide-stencil technique is the critical enabler, not the learning algorithm. SINDy’s advantage is speed (<1<1 s vs. $\sim$835 s); its limitation is the absence of dynamical validation—SINDy returns sparse coefficients but cannot roll out trajectories or verify energy conservation. MAL’s energy-conservation diagnostic provides this additional layer.
@@
(2) HNN [7] and LNN [8]: Trained on identical data (Supplemental Table S6). HNN (17K parameters, 113 s) achieved excellent energy conservation ($\sigma_{H}=4.1\times 10^{-4}$) but learns a black-box Hamiltonian with no interpretable symbolic form. LNN (17K parameters, 242 s) suffered from Hessian singularities, with validation loss diverging to $10^{6}$ throughout training—a known failure mode for LNNs on noisy data where the mass matrix becomes ill-conditioned. Neither method performs basis selection or yields a symbolic force law.
@@
(3) Mathematical LLM (Qwen2-Math 7B): 0% on inverse problems, 61% on forward derivation (Supplemental Material, Section S1).
@@
(4) Physics-informed NN [2]: 85% when the inverse-square law is pre-specified, but cannot identify the law de novo.
@@
Summary of niche. MAL occupies a distinct position: it combines interpretable symbolic basis selection (shared with SINDy [3]) with dynamical rollout and energy-conservation validation (shared with HNN [7]), while adding explicit sparsity-driven efficiency optimization. HNNs/LNNs [7, 8] guarantee conservation but produce black-box energy functions; SINDy [3] yields symbolic expressions but without dynamical consistency checks; MAL provides both within an energy-constrained framework.
III Discussion
III.1 Key findings
Our results establish four principal findings:
(1) Sparsity constraints improve physical law identification. Embedding energy minimization ($E_{\min}$) alongside information maximization ($I_{\max}$) in the triple-action training objective constrains the model selection problem from an unconstrained search to a tractable optimization. The $E_{\mathrm{min}}$ subsystem drives gate crystallization (basis selection) that does not occur under prediction error alone, while reducing training energy by 40%.
@@ III.1 Key findings
(3) Intrinsic crystallization timescales. The gate competition dynamics exhibit universal timescales: crystallization span $\Delta t_{\text{span}}=36.2\pm 4.1$ epochs and geometric growth rate $\gamma=1.137\pm 0.013$ per epoch, independent of which basis wins or the initialization seed. We also observe that major gate transitions coincide with epochs where the schedule ratio $\alpha_{E}/\tau$ passes through integer values, though this may simply reflect the density of rationals along any smooth trajectory through parameter space. Whether the crystallization dynamics are genuinely sensitive to these passages requires perturbation experiments, which we leave to future work.
@@
(4) Architectural sparsification from energy minimization. Clune et al. [11] demonstrated that networks evolving under pressure to minimize connection costs spontaneously develop modular architectures. Our work shows a related pattern: MAL’s sparse gate selection ($C_{\mathrm{gate}}=0.99$) emerges from $\mathcal{L}_{E_{\min}}$ without a pre-programmed bias toward any particular basis. This supports the broader principle that cost-constrained optimization drives architectural sparsification in both biological [12] and artificial systems.
III.2 Broader implications
@@
Energy-efficient scientific machine learning. MAL’s 40% energy reduction over prediction-error-only baselines demonstrates that explicit sparsity constraints can improve computational efficiency without sacrificing task performance. The mechanism is architectural: $\mathcal{L}_{E_{\min}}$ drives early gate crystallization, reducing the effective parameter count and shortening convergence. This principle—that sparsity-inducing regularization simultaneously improves interpretability and efficiency—may generalize to larger-scale scientific discovery tasks.
@@
Noether-based model selection. A persistent challenge in scientific machine learning is distinguishing genuine physical understanding from statistical correlation [16]. The energy-conservation criterion—preferring models that preserve Hamiltonian dynamics over long-horizon rollouts—provides a physics-grounded approach complementary to recent work on learning conserved quantities [9, 10]. Models that violate energy conservation reveal their failure to capture true causal structure, even when matching short-term predictions.
III.3 Limitations and future directions
@@
Fixed basis library. MAL requires pre-specifying candidate force laws $\{\phi_{i}\}$, with the correct answer included. This makes MAL a model selection framework, not an open-ended discovery system. Basis library sensitivity experiments (Supplemental Material, Section S4F) quantify this limitation: adding near-confounders ($r^{-2.5}$, $r^{-1.5}$) reduces the correct-basis rate from 50% to 20%, while removing the correct basis ($r^{-2}$) causes the system to split between alternatives. Distant additions ($r^{2}$, $r^{-4}$, $\ln r$) have no effect. In all cases, the energy-conservation diagnostic remains informative, and uniformly elevated $\sigma_{H}$ across seeds flags potential library inadequacy. Extending to genuinely autonomous discovery—where the algorithm constructs novel functional forms not in the library—requires integration with open-ended symbolic regression methods [4, 5] or LLM-guided symbolic search, which is an important direction for future work.
@@
IV Conclusion
We have demonstrated that minimum-action learning—neural network training guided by a triple-action functional combining information maximization, energy minimization, and symmetry enforcement—enables energy-constrained identification of physical force laws from noisy data. Validated on two benchmarks (Kepler gravity and Hooke’s linear restoring force), the key technical contributions are: (1) the wide-stencil noise reduction technique that transforms an intractable problem (SNR $\sim 0.02$) into a learnable one (SNR $\sim 1.6$); (2) the soft-to-discrete gate sharpening mechanism that achieves basis selection through energy-driven architectural crystallization, with a physics-informed initialization achieving 10/10 correct selection; and (3) the energy-conservation-based model selection criterion that discriminates correct from incorrect physics via long-horizon Hamiltonian conservation. Direct comparison against SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: interpretable, energy-constrained model selection combining symbolic basis identification with dynamical validation.
Future work will extend MAL to non-central forces, dissipative systems, higher-dimensional problems, and real observational data (e.g., planetary ephemerides), and will test whether the intrinsic crystallization timescales and schedule-geometry coincidences observed here generalize across problem classes.
V Methods
V.1 Data generation
@@ V.3 Loss function
V.4 Training protocol
Two-phase schedule:
@@ V.5 Validation metrics
V.6 Code and data availability